OpenStudy (anonymous):

use the given base to write the number as an exponential expression 216 (base 6) 216 (Base6)=

OpenStudy (whpalmer4):

How many 6's do you have to multiply together to get 216?

OpenStudy (anonymous):

36

OpenStudy (anonymous):

@whpalmer4 36

OpenStudy (whpalmer4):

no, that's not quite the same, that's adding 36 6's together. You multiply together 3 2's to get 8: 2*2*2 = 8 How many 6's to get 216?

OpenStudy (anonymous):

3 @whpalmer4 6*6*6=216

OpenStudy (whpalmer4):

Yes, that's right. So what is the answer to the problem? Write 6*6*6 = 216 as an exponential expression...

OpenStudy (anonymous):

thanks

OpenStudy (whpalmer4):

What's your answer?

OpenStudy (anonymous):

6 to the3rd power 

OpenStudy (whpalmer4):

That's correct!